OpenStudy (anonymous):

At the Octeplex Cinema, the Smith Family spends $29 to buy 2 adult tickets and 4 children tickets. At the same cinema, The Johnson family spends $37 to buy 4 adult tickets and 2 children tickets. What is the price of an adult ticket? $3.50 $4 $7.50 $8 $9.50

OpenStudy (anonymous):

system of equations: x:adults y:children 2x+4y=29 4x+2y=37 solve to find x and y x is your answer

OpenStudy (anonymous):

x= 15/2 ?

OpenStudy (anonymous):

so it's 7.50

OpenStudy (anonymous):

yes